A room (10 m× 5 m × 5 m)contains moist air at 38°C / 1 bar with relative humidity 70%. Determine the specific humidity?(where, Corresponding to the dry bulb temperature of 38° C, the saturation pressure of water vapour Ps=0.06624, From the Steam Table)



1
Expert's answer
2021-06-28T08:53:00-0400

V = "10m \u00d7 5m \u00d7 5m" = 250m³

density of air = 1.225 kg/m³

mass of air = 1.225kg/m³ × 250m³ = 306.25 kg


water = 83 g/m³

70% humidity = 70% × 83g/m³ = 0.0581kg/m³

mass of water = 0.0581 × 250 = 14.525kg


Specific Humidity = mass of water vapor per unit mass of air = 14.525/306.25 = 0.047